Please help me with solving:
2x-y=1
4x-2y=2

Found 2 solutions by mananth, richard1234:
Answer by mananth(16946)

2x-y=1
y=2x-1

4x-2y=2
substitute y
4x-2(2x-1)=2
4x-4x+2=2
2=2
meaning that the lines are co incident and have innumerable solutions
.

Answer by richard1234(7193)

The second equation is obtained by multiplying the first equation by 2, hence they are equivalent. Infinitely many solutions